A violin string of length 44 cm and mass 1.1 g has a frequency of 534 Hz when it is vibrating in its fundamental mode.

(a) What is the wavelength of the standing wave on the string?
_______ cm

(b) What is the tension in the string?
_______N

(c) Where should you place your finger to increase the frequency to 684 Hz?
__________ cm from the fixed end of the string (from the peg of the violin)
Answer #1

a] For the fundamental mode, two nodes are set up at the edges and one antinode in the middle.

This implies that the wavelength in such a case is twice the length of the string

=> \small \lambda = 2L = 88 cm=0.88m

b] Mass per unit length of the string = m/L = 1.1 x 10-3/0.44 = 2.5 x 10-3 kg/m

and the speed of the wave will be:

v = wavelength x frequency = 469.92 Hz

speed of the wave is given by:

\small v = \sqrt{\frac{T}{m/L}}

so, \small (469.92)^2 = \frac{T}{2.5\times 10^{-3}}

=> T = 552.06 N.

c]

Use the relation

\small \frac{f_2}{f_1}=\frac{L_1}{L_2}

=> \small \frac{684}{534}=\frac{44}{L_2}

=> L2 = 34.35 cm

so, this effective length can be created by plucking the string at a distance of D = L1 - L2 = 9.65 cm from the fixed end of the string.